Let vector a and b be two non-null vectors such that |vector a +b| =|vector a-2b| then the values of |vector a| / |vector b| may be:???
 

Grade:11

1 Answers

ADITI SINGH
7 years ago
oh i am not sure but see what i did

//bold alphabets are vectors
+ 2is parallet to c
so 

c = k(a + 2b)  lets call it equation 1 as we will use it later
 // where k is some scalar

also b + 2is paraller to a

so  b + 2c  = m a // for some scalar m 

now put  value of c from equation 1


b + 2ka + 4kb = ma

//rearrranging gives

(4k + 1)b = (m - 2k )a

now we can have 4k + 1 = 0

gives k = -1/4

so put this value of k in equation 1

 we get c = (-1/4)(a + 2b)


which gives a + 2b = -4c

which is a + 2b + 4b = 0

see if my method is wrong somewher